OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

paralell lines have same slope so slope=3 and now use this .. (y-y1)=m(x-x1) plug in values and solve..m=slope

OpenStudy (anonymous):

whats the y- y1 and x-x1 how do you know the difference.... sorry im not very good at math lol

OpenStudy (anonymous):

here you have (x1,y1)=(4,-2) put these vaues at thier place and solve..

OpenStudy (anonymous):

(y-4)=3(x+2) -4y=3x+6 do i have it right so far?